OpenStudy (anonymous):

slope intercept form for the line that goes through (-6,-2) and is parallel to -7x+3

OpenStudy (anonymous):

I need to find an equation in slope intercept form

OpenStudy (katrinakaif):

use the equation (y - y1) = m (x-x1) SInce its parallel, the slope has to be the same as the equation given (m = -7) NOw plug in (-6,-2) as (x1,y1) in the equation concludes (y+2) = -7(x +6) and solve

OpenStudy (anonymous):

Okay now how do I solve it? Do I add them?

OpenStudy (katrinakaif):

(y + 2) = -7x -42 y = -7x -44

OpenStudy (anonymous):

-7x+3...is it not equated to other no.?? is there no equal sign?

OpenStudy (katrinakaif):

its presumed.

OpenStudy (katrinakaif):

Since it is a slope-intercept form, it has to be equated to "y" concluding y = -7x + 3

OpenStudy (anonymous):

Okay thank you so much

OpenStudy (katrinakaif):

Anytime sweetie =)